# Can one hear the full nonlinearity of a PDE from its small excitations?

### v1

@@ 2 Principal result: the inverse linearization protocol for PDEs of the sine-Gordon type
We claim the following:
@@
The expansion (8) is substituted to the nonlinear equation (3), and subsequently, only the terms of the first order in $\delta u(x)$ are kept. This procedure results in the following eigenmode equation:
@@ 3 Justification of the inverse linearization procedure (4), (5), (6)
First of all, in Section 2, we assumed that the inverse $\chi(\bar{u})$ of $\bar{u}(x)$ (see (4)) exists. Thanks to the positivity of $\delta u_{0}(x)$, the existence of this inverse is however guaranteed: $\bar{u}(x)$ is a monotonic function of $x$.
@@
However, the proof above tell one nothing about why the recipe works. The cornerstone for understanding the physics behind the connection formulas (4,5, 6) is the Goldstone theorem: if a PDE possesses a continuous symmetry, the action of the generator of this symmetry on a stationary solution is a small excitation around this solution, with a zero frequency. In our case, the symmetry in question is the translational invariance of the PDE (3) in question. Therefore the spatial derivative of a stationary solution is a legitimate zero-frequency mode of the linear stability analysis equation around that stationary solution. Conversely, the kink is proportional to an integral over that eigenstate; hence the connection (4), where it is assumed that it is the ground state of (1) that corresponds to the Golstone mode. This is the connection between a nonlinear PDE and its linearization that we are exploiting in this work. This connection can also be seen explicitly by differentiating (7) with respect to $x$ and observing that the result is identical to the linearization (9) with $\omega=0$, applied to the kink state $\bar{u}$.
@@ 4 Worked examples
First of all, we would like to verify that our procedure returns the sine-Gordon equation, if applied to its linearization around its stationary soliton. In that case the potential of the linearized problem is the Pöschl-Teller potential [4], $V(x)=-n(n+1)\operatorname{sech}^{2}(x)$, with $n=1$ (see row 1 of Table 1). Using its ground state, $\delta u_{0}=\operatorname{sech}(x)$, to initialize the protocol in the protocol (4), (5), (6), with $A=2$ and $B=0$ we indeed recover the sine-Gordon equation.
@@
The $\delta$-well and the harmonic oscillator case are also included in our study (rows 5 and 6 respectively). In the former case, the nonlinearity $F(u)$ is a piece-wise-linear, periodic function of $u$: we named the resulting PDE a sawtooth-Gordon equation [5]. In the latter case, a function $F(u)$ that is expressed in terms of the inverse of the error function emerges. Similarly to the $n=4$ Pöschl-Teller case, harmonic oscillator leads to a nonanalytic PDE.
@@ 5 Conclusions and outlook
Next, in [5], it has been shown that integrable PDEs produce reflectionless linear stability analysis equations. Regretfully, the converse is not true, as witnessed by the the row 2 of Table 1 pertaining to formed the $n=2$ Pöschl-Teller Hamiltonian (see [6] on its transparency) and a particular $\phi^{4}$ model ($F(u)=-2u+2u^{3}$) whose non-integrable behavior has been well documented [7].
Looking ahead, several directions future research can be foreseen. First note that the ground state of the linear Schrödinger equation (1) is not the only eigenstate that produces a monotonic (and hence invertible) kink state $\bar{u}(x)$ (see (4)). For $V(x)$ supporting a single bound state or no bound states at all, the $E=0$ eigenstate can be chosen to be monotonic, leading to a monotonic $\bar{u}(x)$. In that case the equation (18) needs to be modifies as
@@
As an example of an inverse linearization pertaining to the Case 2, one may consider the fist excited state $\delta u_{1}(x)$ (of energy $E_{1}=-1$) of the $n=2$ Pöschl-Teller problem leading to a different kind of the $\phi^{4}$ model, with $F(u)=u-2u^{3}$. The analogue of the equation (18) will read
@@
We were not able to identify any instances corresponding to the Case 3.
It should be noted that whenever the Goldstone mode is not represented by the ground state of the linearization problem, the corresponding kink soliton is dynamically unstable. Indeed, in that case, the ground state, as well as any other eigensate with energy below the Goldstone value will have an imaginary frequency.

### v2

@@ 2 Principal result: the inverse linearization protocol for PDEs of the sine-Gordon type
It is assumed that $F(u)$ is smooth. We claim the following:
@@
The expansion (8) is substituted in the nonlinear equation (3), and subsequently, only the terms of the first order in $\delta u(x)$ are kept. This procedure results in the following eigenmode equation:
@@ 3 Justification of the inverse linearization procedure (4), (5), (6)
First of all, in Section 2, we assumed that the inverse $\chi(\bar{u})$ of $\bar{u}(x)$ (see (4)) exists. Thanks to the positivity of $\delta u_{0}(x)$, the existence of this inverse is, however, guaranteed: $\bar{u}(x)$ is a monotonic function of $x$.
@@
However, the proof above tells one nothing about why the recipe works. The cornerstone for understanding the physics behind the connection formulas (4,5, 6) is the Goldstone theorem: if a PDE possesses a continuous symmetry, the action of the generator of this symmetry on a stationary solution is a small excitation around this solution, with a zero frequency. In our case, the symmetry in question is the translational invariance of the PDE (3) in question. Therefore the spatial derivative of a stationary solution is a legitimate zero-frequency mode of the linear stability analysis equation around that stationary solution. Conversely, the kink is proportional to an integral over that eigenstate; hence the connection (4), where it is assumed that it is the ground state of (1) that corresponds to the Goldstone mode. This is the connection between a nonlinear PDE and its linearization that we are exploiting in this work. This connection can also be seen explicitly by differentiating (7) with respect to $x$ and observing that the result is identical to the linearization (9) with $\omega=0$, applied to the kink state $\bar{u}$.
@@ 4 Worked examples
First of all, we would like to verify that our procedure returns the sine-Gordon equation, if applied to its linearization around its stationary soliton. In that case the potential of the linearized problem is the Pöschl-Teller potential [4], $V(x)=-n(n+1)\operatorname{sech}^{2}(x)$, with $n=1$ (see row 1 of Table 1). Using its ground state, $\delta u_{0}=\operatorname{sech}(x)$, to initialize the protocol in the protocol (4), (5), (6), we obtain, by applying (4):
$\bar{u}(x)=A(B+2\arctan{e^{x}})\,\,$ (16)
Define the function:
$u(x)=\frac{A}{2}\sin(\frac{2\bar{u}(x)}{A}-2B)$ (17)
Then following (5) and (6), the nonlinearity in the blank-Gordon equation $F(\bar{u}))$ is given by:
$F({\bar{u}})=\frac{A}{2}\sin(\frac{2}{A}\bar{u}-2B)=\sin(u)$ (18)
For $A=2,B=0$ we recover sine-Gordon equation. For arbitrary $A$ and $B$, the linear change of variables $\bar{u}\to u$, given by (17) yields the sine-Gordon equation in $u$, so the constants $A$ and $B$ do not affect the dynamics.
In fact the constants $A$ and $B$ do not affect the dynamics, as solutions that differ only through the choice of $A$ and $B$ are related through an affine transformation. Suppose $u(x)$ is a stationary solution to the blank-Gordon equation with nonlinearity $F(u)$. That is $u^{\prime\prime}(x)=F(u)$. Suppose that:
$v(x)=A(u(x)+B)$ $u(x)=\frac{v(x)-B}{A}$ (19)
Both $u(x),v(x)$ can be obtained from (4), corresponding to an arbitrary choice of $A,B$.
Then:
$v^{\prime\prime}=\bar{F}(v):=AF(\frac{v(x)-B}{A})+B$
Then changing variables from $v\to u$ yields the equation $u^{\prime\prime}=F(u)$, so that the dynamics are the same regardless of the choice of A and B. Since $A,B$ do not affect the dynamics, the examples shown in the table use values $A,B$ that yield well-known or particularly simple $F(u)$.
@@
The $\delta$-well and the harmonic oscillator case are also included in our study (rows 5 and 6 respectively). In the former case, the nonlinearity $F(u)$ is a piece-wise-linear, periodic function of $u$: we named the resulting PDE a sawtooth-Gordon equation [5]. This form of $F(u)$ runs against our assumption of smooth nonlinearity, but our scheme still holds in the sense of distributions in this case. In the latter case, a function $F(u)$ that is expressed in terms of the inverse of the error function emerges. Similarly to the $n=4$ Pöschl-Teller case, harmonic oscillator leads to a nonanalytic PDE.
@@ 5 Conclusions and outlook
Next, in [5], it has been shown that integrable PDEs produce reflectionless linear stability analysis equations. Regretfully, the converse is not true, as witnessed by row 2 of Table 1 pertaining to the $n=2$ Pöschl-Teller Hamiltonian (see [6] on its transparency) and a particular $\phi^{4}$ model ($F(u)=-2u+2u^{3}$) whose non-integrable behavior has been well documented [7].
Looking ahead, several directions of future research can be foreseen. First note that the ground state of the linear Schrödinger equation (1) is not the only eigenstate that produces a monotonic (and hence invertible) kink state $\bar{u}(x)$ (see (4)). For $V(x)$ supporting a single bound state or no bound states at all, the $E=0$ eigenstate can be chosen to be monotonic, leading to a monotonic $\bar{u}(x)$. In that case the equation (22) needs to be modified as
@@
As an example of an inverse linearization pertaining to the Case 2, one may consider the first excited state $\delta u_{1}(x)$ (of energy $E_{1}=-1$) of the $n=2$ Pöschl-Teller problem leading to a different kind of the $\phi^{4}$ model, with $F(u)=u-2u^{3}$. The analogue of the equation (22) will read
@@
We were not able to identify any instances corresponding to Case 3.
It should be noted that whenever the Goldstone mode is not represented by the ground state of the linearization problem, the corresponding kink soliton is dynamically unstable. Indeed, in that case, the ground state, as well as any other eigenstate with energy below the Goldstone value, will have an imaginary frequency.